OpenStudy (pumpkin1102):

Joy and Cullen put some money into their money boxes every week. The amount of money (y), in dollars, in their money boxes after a certain amount of time (x), in weeks, is shown by the equations below: Joy: y = 30x + 20 Cullen: y = 20x + 90 After how many weeks will Joy and Cullen have the same amount of money in their money boxes, and what is that amount? 6 weeks, $150 230 weeks, $7 7 weeks, $150 7 weeks, $230

OpenStudy (error1603):

its not A and C

OpenStudy (pumpkin1102):

I think it is B

OpenStudy (error1603):

y = 30x + 20 20x + 90 = 30x + 20 ... Replace y with 20x+90 20x + 90 - 90 = 30x + 20 - 90 ... Subtract 90 from both sides. 20x + 0 = 30x - 70 20x = 30x - 70 20x = -70 + 30x 20x - 30x = -70 + 30x - 30x ... Subtract 30x from both sides. -10x = -70 + 0x -10x = -70 x = ?

OpenStudy (pumpkin1102):

urf idk

OpenStudy (error1603):

urf?

OpenStudy (pumpkin1102):

urf just meaning gosh idk

OpenStudy (error1603):

oh x=7

OpenStudy (pumpkin1102):

so it is D?

OpenStudy (error1603):

Yep

OpenStudy (pumpkin1102):

thx

OpenStudy (error1603):

No problem